Shift report 20020102 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 02 2002 - 00:03:27 EST

  • Next message: Apache: "Shift report 20020102 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Jan  2 00:03:27 2002, Shift: 16:00-24:00
    Supervisor: fv, with: (sjs,hiro) parttime
    Log book#11 pages 130 - 134
    
    Shift Summary:
    ------------------------------------------------------------
    Much time spend on machine injection.
    store one  17:30 - 19 inel rate ~7K
    and ongoing started 22:35 inel rate 3.5K.
    
    Runs FS 12 deg 1/8A 6234, 6235, 6236(in progress)
         45K trigger 6 
    MRS 90 deg 2A ~40K trigger 3.
    
    Checked timing for inel counters. Made small adjustments for
    R1,L1. They are now in trigger (For runs 6227-6233) the selection of trigger counters was L2-4 and R1-3 (not R2-4).
    
    details:
    
    
    16:30 One unsuccesfull ramp with too little yellow beam to keep.
    
    17:30 	Succesfull ramp.
    	
    17:35	Start run 6234 
    	FS 1/8A    
    	MRS 90 deg 2A
    
    At 18:05 some machine activity caused the background rates 
    to change considerably (*2 in H1,H2) 
    See attached pict in logbook page 131.
    
    18:30 	Stop run 6234.
    
    18:30-19:30 Look and investigate inel trigger,
    	checking delays settings, and changing OR to R1-R4 and
    	L1-L4. Rtaes only went up ~ 5500-> 6500
    19:35 	Run 6235 started.
    
    19:45 paused run 6235 for pol measurements.
    
    19:50	RHIC quench link interlock
    
    22:25	By reading the RHIC e-log saw the plot + comment that the
    	quenchlink relates to the Phobos magnet. The other interesting
    	BAD item is that the increase noticed in background rates at 8:05-815
    	correlates with a first turn on of the Phobos magnet!!
    
    22:30 	The next fill done. Looks promising, but not very high.
    	~3.8 K (with R1-4 AND L1-4 in inel trigger)
    
    22:35	Start run 6236
    	Actual polarization measurement did not take place 
    	until 23:03-23:05 (This may have been phobos magnet again)
    
    23:25	It is still called machine setup - presumably MCR is
    	doing pol 
    
    	
    
    
    
    
    
    Beam so far
    
    17:30-19:45 == 2h 15 min
    22:30-         1H 30 min  (3:45)
    
    
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Jan 02 2002 - 00:03:44 EST